OpenStudy (agentnao):

What is the focus of the parabola y = –1/16 (x + 3)^2 – 6 ?

OpenStudy (amistre64):

its the point that is on the axis of symmetry a distance of "a" from the vertex :)

OpenStudy (amistre64):

the vertex looks to be (-3,-6)

OpenStudy (amistre64):

(y+6) = –1/16 (x + 3)^2 -16 (y+6) = (x+3)^2 relate this to: 4a y = x^2 4a = -16; a = 4

OpenStudy (amistre64):

F = (1,6) maybe?

OpenStudy (amistre64):

ugh, (-3,-10) according to the wolf

OpenStudy (agentnao):

Hm, well, my notes say that the focus is (0, c) (which is 4) but the options are ...

OpenStudy (agentnao):

Oh. Lol, yes, (-3, -10) but where did they get the 10 from?

OpenStudy (amistre64):

i see what I did :) the vertex is at (-3,-6) the focus is 4 away in the open direction; the open direction is down (-3,-6-4) = (-3,-10)

OpenStudy (amistre64):



OpenStudy (agentnao):

Ohh. Oh my gosh, that makes so much sense! And if it were a sideways parabola, then "a" would have been added to the x coordinate.

OpenStudy (amistre64):

yep

OpenStudy (agentnao):

Thank you so much :) I'll def remember that!

OpenStudy (amistre64):

ill prolly forget it again lol

OpenStudy (agentnao):

Haha, well, I suppose eventually I will too, but at least I'll know it for the test :)